Find the indicated probabilities using the geometric distribution, the Poisson distribution, or the binomial distribution. Then determine if the events are unusual. If convenient, use the appropriate probability table or technology to find the probabilities.
A major hurricane is a hurricane with wind speeds of 111 miles per hour or greater. During the last century, the mean number of major hurricanes to strike a certain country's mainland per year was about
0.65.
Find the probability that in a given year (a) exactly one major hurricane will strike the mainland, (b) at most one major hurricane will strike the mainland, and (c) more than one major hurricane will strike the mainland.

Answer by Boreal(15235)

This is a Poisson--not a fixed number of trials (excludes binomial), proportional to time, low probability, and theoretically can be infinite
for exactly 1, lambda is 0.65 and probability is e^(-0.65) 0.65^1/1!=0.3393
for 0, it is e^(-0.65)=0.5220
so at most 1 hurricane has probability of the sum of those two or 0.8613
more than 1 hurricane would be the complement or 1-0.8613=0.1387 probability.